An object traveling along the x axis at constant acceleration has a velocity of +10 m/s when it is at x = 6.0 m and of +15 m/s when it is at x = 10.0 m. What is its acceleration? 


1
Expert's answer
2021-02-25T11:38:53-0500

The acceleration of the object

"a=\frac{v_2^2-v_1^2}{2(x_2-x_1)}\\
=\frac{15^2-10^2}{2*(10-6.0)}=15.6\:\rm m/s^2"